Complete each probability rule. If E and F are mutually exclusive events then P(E UF)- P(E)+PF) P(E) P(E I F) PE) + P(F) # 1 P(E)- 1- P(E) 1-P(E) If E and F are mutually exclusive events then P(E)+ P(F) If E and F are independent events, then P(E IF) number of outcomes in E number of outcomes in sample spaceE)PF
Answer #1

if E and F are mutually exclusive then P(E U F)=P(E)+P(F)

P(Ec) =1-P(E)

if E and F are mutually exclusive then P(E n F)=0

if E and F are independent events then P(E n F)=P(E)*P(F)

number of outcomes in E/number of outcomes in sample space=P(E)